OpenStudy (anonymous):

Can someone write (1,-8); m=5 in standard form and point slope form?

OpenStudy (lgbasallote):

first..use the point slope form \[y - y_1 = m(x-x_1)\] in this case your is 5 and x1 is 1 and y1 is -8 so the equation becomes \[y -(-8) = 5(x-1)\] following so far?

OpenStudy (anonymous):

Yes I am

OpenStudy (lgbasallote):

so that is the point-slope form..to get the standard from, just simplify \[\implies y + 8 = 5x - 5\] put everything on one side \[\implies 5x - y - 5 - 8 = 0\] combine like terms... \[\implies 5x - y - 13 = 0\] ^standard form got it?